OpenStudy (fanduekisses):

Find the derivative of f(x) = -12x2 + 9x at x = 6

OpenStudy (fanduekisses):

Do I plug in for x

OpenStudy (anonymous):

d/dx = -24 x + 9 Plug in x = 6 you got -135

OpenStudy (anonymous):

Plug in after you take the derivative

OpenStudy (fanduekisses):

how do I take the derivative?

OpenStudy (anonymous):

\[\frac{ d }{ dx } f(x) = n * x ^{n-1}\]. That's the simplest definition. In your cause move 2 down and multiply with -12. then on the exponential part subtract 1 from 2. You have (2) (-12) x^ (2-1), which equates to -24x

OpenStudy (anonymous):

as for 9x, think of it as 9x^1. So (1)(9) x^(1-0) = 9

OpenStudy (fanduekisses):

thank you